OpenStudy (anonymous):

In a triangle, the middle length side is 3 more than the shortest side, and the longest side is double the middle length side. The perimeter of the triangle is 45 cm. Find the lengths of the sides.

OpenStudy (wikiemol):

you will have a triangle that looks like this:  its perimeter is 45 so now you have the equation x + 3x + 6x = 45

jimthompson5910 (jim_thompson5910):

3 more doesn't mean 3x, it means x+3

jimthompson5910 (jim_thompson5910):

let x = shortest side y = middle side z = longest side the middle length side is 3 more than the shortest side so y = x+3 the longest side is double the middle length side which means z = 2y z = 2(x+3) z = 2x + 6 ------------------------------------------------ All sides must add to 45, so x+y+z = 45 x+x+3+2x+6 = 45 I'll let you finish up

OpenStudy (wikiemol):

Right! Sorry...  and x + x+3 +2(x+3) = 45

OpenStudy (anonymous):

so the middle side would be 13?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

Then what would it be?

OpenStudy (wikiemol):

I got x = 9

jimthompson5910 (jim_thompson5910):

first solve for x x+x+3+2x+6=45 4x+9=45 4x=45-9 4x=36 x=(36)/(4) x=9

jimthompson5910 (jim_thompson5910):

then use this to find y y = x+3 y = 9+3 y = 12 y was defined to be the middle side, so the middle is 12

OpenStudy (anonymous):

i got 10. okay i see i used 49 instead of 45....

jimthompson5910 (jim_thompson5910):

ah that's the culprit

OpenStudy (anonymous):

Alright, i get it now.

jimthompson5910 (jim_thompson5910):

ok that's great